OpenStudy (anonymous):

Factor the following polynomial completely. -x 2 y 2 + x 4 + 4y 2 - 4x 2

OpenStudy (anonymous):

I Don't Get This At All

OpenStudy (ikram002p):

do you mean ? \(-x ^2 y ^2 + x ^4 + 4y^ 2 - 4x^ 2\)

OpenStudy (anonymous):

Yess

OpenStudy (ikram002p):

\(-x ^2 y ^2 + x ^4 + 4y^ 2 - 4x^ 2 =-y^2(x^2-4)+ x^2(x^2-4)=(x^2-4)(x^2-y^2)\)

OpenStudy (ikram002p):

\((x^2-4)(x^2-y^2)=(x-2)(x+2)(x-y)(x+y)\)

OpenStudy (ikram002p):

done ^

OpenStudy (anonymous):

My Choices are A).(x - 2)(x - 2)(x + y)(x - y) B).(x + 2)(x + 2)(x + y)(x - y) C).(x + 2)(x - 2)(x + y)(x - y)

OpenStudy (ikram002p):

ok lol i solved it for u, what do u think which one is solution ?

OpenStudy (ikram002p):

check C ^

OpenStudy (anonymous):

Aweee I Feel Slow

OpenStudy (ikram002p):

haha its ok :P it happens , me too feels tierd